OpenStudy (anonymous):

Rearrange 0=100(200π)cos(200πt+π/4) to make t the subject.

OpenStudy (amistre64):

divide, arc, subtract, divide atain

OpenStudy (amistre64):

*again

OpenStudy (loser66):

show me how to do please

OpenStudy (amistre64):

i just did lol

OpenStudy (amistre64):

\[0 = k~trig(pt+n)\] \[0 =trig(pt+n)\] \[arc(0) =pt+n\] \[arc(0)-n =pt\] \[\frac 1p(arc(0)-n) =t\]

OpenStudy (loser66):

at the end, "solve for t" is what you mean, right?

OpenStudy (amistre64):

no, not solve for t, just algebra it into an equation for t

OpenStudy (amistre64):

to be fair, arccos(0) = pi/2 but any odd multiple of pi/2 is sufficient

OpenStudy (anonymous):

thank you @amistre64, big help. I was making a mistake when it came to the cos, not using arc!! thank you again.

OpenStudy (amistre64):

your welcome

OpenStudy (loser66):

isolate t? right?@amistre64

OpenStudy (amistre64):

yes

OpenStudy (loser66):

ok. Got it. thanks for explanation

OpenStudy (amistre64):

yep